# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ 2.6. Directional Expansiveness for $\mathbb{R}^{d}$
It turns out that the correct definition of direction ${\mathcal{V}}$ expansiveness for an $\mathbb{R}^{d}$ action $T$ is not simply the definition of expansiveness for $T$, restricted to ${\mathcal{V}}$, even though this does work in some directional theories. The restriction of $T$ to ${\mathcal{V}}$, defined $(T|_{\mathcal{V}})^{\boldsymbol{t}}:=T^{\boldsymbol{t}}$ for ${\boldsymbol{t}}\in{\mathcal{V}}$, is a ${\mathcal{V}}$ action, but can be made into an $\mathbb{R}^{e}$ by a choice of a basis (see e.g., [3]). This does not affect questions of expansiveness.
@@ 3. Tiling dynamical systems
In this section we define finite local complexity (FLC) tiling spaces and tiling dynamical systems, and describe their basic properties (see e.g., [28] or [1]) For simplicity, we will always assume $d=2$.
@@ 3.1. Basic definitions
A tile is a subset $D\subseteq\mathbb{R}^{2}$ homeomorphic to a closed disk, with a Lebesgue measure zero boundary (e.g., a convex polygon). A tiling $x$ of $\mathbb{R}^{2}$ is a set of the tiles that cover $\mathbb{R}^{2}$, intersecting only on their boundaries. A tiling patch $y$ in $\mathbb{R}^{2}$ is a finite collection of the tiles whose union is a closed topological disc, and a tiling swatch $y$ is a collection of tiles whose union is closed and simply connected.
If $x$ is a tiling of $\mathbb{R}^{2}$, and $S\subseteq\mathbb{R}^{2}$, we let $x[S]$ be the smallest patch or swatch, the union of whose tiles contains $S$. Two tiles, patches, swatches or tilings are said to be equivalent if they are congruent via a translation. Sometimes geometrically equivalent tiles are “decorated” with colors or markings that make them inequivalent. We denote by $T^{{\boldsymbol{t}}}x:=\{D-{{\boldsymbol{t}}}:D\in x\}$ the translation of a tiling (or tile or patch or swatch) $x$ by ${{\boldsymbol{t}}}\in\mathbb{R}^{2}$. We call a set $p$ of inequivalent tiles in $\mathbb{R}^{2}$ a prototile set. A tiling (or path or swatch) $x$ by $p$ means $x$ each tile in $x$ is equivalent to a tile in $p$. We want to look at sets of all tilings by $p$, but without further restrictions, such a set is often too large.
@@
There is a well known compact metric topology on any FLC tiling spaces $X_{p}$ in which two tilings are close if, after a small translation, they agree on a large ball around the origin. Here is one version of such a metric $\rho$ (see [28], [34]). Let $\Delta_{p}:=\min_{D\in p}{\rm diam}(D)$. For $x,y\in X_{p}$, $x\neq y$, let
$R(x,y):=\sup\left\{r:(T^{{\boldsymbol{r}}_{x}}x)[B_{r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{r}}_{y}}y)[B_{r}({{\boldsymbol{0}}})]\text{ some }||{\boldsymbol{r}}_{x}||,||{\boldsymbol{r}}_{y}||<1/r\right\},$
and define the tiling metric $\rho(x,y):=\min(\Delta_{p},1/R(x,y))$. A well known result of Rudolph [33] is that an FLC full tiling space $X_{p}$ is a compact metric space (see also [28], [34]). The following will be useful.
@@
    If $\rho(x,y)<\delta<\min(1/3,\Delta_{p})$ and $r=1/\delta$, then there exists ${\boldsymbol{t}}\in\mathbb{R}^{d}$ with $||{\boldsymbol{t}}||<\delta$ so that $x[B_{(8/9)r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{t}}}y)[B_{(8/9)r}({{\boldsymbol{0}}})]$.
@@
An FLC tiling space in general is any closed $T$-invariant subset $X\subseteq X_{p}$ where $X_{p}$ is a full tiling space. We call $(X,\mathbb{R}^{2},T)$ an FLC tiling dynamical system.
Remark 3.4.
Let $X$ and $X^{\prime}$ be tiling spaces. For $R>0$, consider the sets of patches $X_{\{R\}}:=\{x[B_{R}({\boldsymbol{0}})]:x\in X\}$ and $X^{\prime}_{\{0\}}:=\{x^{\prime}[\{{\boldsymbol{0}}\}]:x^{\prime}\in X^{\prime}\}$. A map $\theta:X\to X^{\prime}$ is called local (see [28]) if there is a map $\Theta:X_{\{R\}}\to X^{\prime}_{\{0\}}$ so that $\theta(x)[\{{\boldsymbol{0}}\}]=\Theta(x[B_{R}({\boldsymbol{0}})])$ for all $x\in X$. Note that a local map is uniformly continuous, and satisfies $\theta(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}\theta(x)$ for $x\in X$ and ${\boldsymbol{t}}\in\mathbb{R}^{2}$. A local map is a factor map if it is onto. If a local homeomorphism $\theta$ has a local inverse, we say $\theta$ is a mutually locally derivable (MLD) topological conjugacy. Local maps between FLC tiling spaces are analogous to sliding block codes between $\mathbb{Z}^{d}$ subshifts. In the subshift case, the Curtis/Lyndon/Hedlund Theorem says every topological conjugacy is a sliding block code. However, for FLC tiling dynamical systems there are topological conjugacies that are not MLD (see [26], [15], [1] and Theorem 6.11).
@@
For $j=0,1,\dots,4$, let ${{\boldsymbol{v}}}_{j}:=({\rm Re}(e^{2\pi ij/5}),{\rm Im}(e^{2\pi ij/5}))$, the $j$th $5$th root of unity, and let ${{\boldsymbol{v}}}^{\prime}_{j}:=(2/5){{\boldsymbol{v}}}_{j}$. We dente the prototile set $p_{5}:=\{{{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:0\leq i<j\leq 4\}$, where ${{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:=\{t_{0}{{\boldsymbol{v}}}^{\prime}_{i}+t_{1}{{\boldsymbol{v}}}^{\prime}_{j}:t_{0},t_{1}\in[0,1]\}$. Then $p_{5}$ consists of $10$ rhombic tiles with edge length $2/5$, five with acute angle $52\pi/$ 5 and five with acute angle $102\pi/$ 10. Figure 1 (a) shows ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{4}$ and ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{3}$. All other tiles in $p_{5}$ are rotations of one of these. We take local rule $p^{(2)}_{5}$ to be the requirement that all tiles meet edge to edge. The corresponding FLC full tiling space satisfies $X_{p_{5}}\neq\emptyset$ since it contains periodic tilings (see Figure 1, (b)).
@@ 3.2. Expansiveness and directional expansiveness for tiling dynamical systems
Given $\epsilon>0$, let $\delta<\min\left(\epsilon,1/3,\Delta_{p}/2\right)$, $r=1/\delta$. Let $B_{0}=B_{8r/9}({{\boldsymbol{0}}})$. For a sequence ${\boldsymbol{t}}_{n}\in\mathbb{R}^{2}$, $n=1,2,\dots$, let $B_{n}:=B_{8r/9}(-{\boldsymbol{t}}_{n})$. Choose the ${\boldsymbol{t}}_{n}$ so that $\bigcup_{n\geq 1}B_{n}=\mathbb{R}^{2}$, and so that $I_{n}:=B_{n+1}\cap B_{n}\neq\emptyset$. Now suppose $x,y\in X_{p}$ satisfy $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)<\delta$ for all ${\boldsymbol{t}}\in\mathbb{R}^{2}$. Then for each $n=1,2,\dots$, $\rho(T^{{\boldsymbol{t}}_{n}}x,T^{{\boldsymbol{t}}_{n}}y)<\delta$, so that by (2) of Lemma 3.2, there exists $||{\boldsymbol{r}}_{n}||<\delta$ so that $(T^{{\boldsymbol{t}}_{n}}x)[B_{0}]=(T^{{\boldsymbol{t}}_{n}+{\boldsymbol{r}}_{n}}y)[B_{0}]$. Equivalently, $x[B_{n}]=(T^{{\boldsymbol{r}}_{n}}y)[B_{n}]$. We claim ${\boldsymbol{r}}_{n}={\boldsymbol{r}}_{1}$ for all $n\in\mathbb{N}$. Thus $x[B_{n}]=(T^{{\boldsymbol{r}}_{1}}y)[B_{n}]$ for all $n$, and since the $B_{n}$ cover $\mathbb{R}^{2}$, $y=T^{{\boldsymbol{r}}_{1}}x$. By (3) of Lemma 3.2, $||{\boldsymbol{r}}_{1}||=\rho(x,y)<\delta<\epsilon$. To prove the claim, we prove ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$ for all $n$. Let $C_{n}:=B_{n}\cap B_{n+1}$. Then $(T^{{\boldsymbol{r}}_{n}}y)[C_{n}]=x[C_{n}]=(T^{{\boldsymbol{r}}_{n+1}}y)[C_{n}]$. Since $||{\boldsymbol{r}}_{n+1}-{\boldsymbol{r}}_{n}||<2\delta<\Delta_{p}$, (1) of Lemma 3.2 implies ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$. ∎
@@
For the induction step, we assume that, for some $k\geq 1$, no ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<rk$ has $||h({\boldsymbol{t}})-{\boldsymbol{t}}||\geq 2\delta/3$, and show the same for ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<(k+1)r$. Suppose to the contrary that some such ${\boldsymbol{t}}={\boldsymbol{t}}_{0}$ satisfies $||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=2\delta/3$. Let ${\boldsymbol{s}}_{0}=kt/(k+1)$ so that ${\boldsymbol{s}}_{0}\in{\mathcal{V}}$, $||{\boldsymbol{s}}_{0}||<kr$ and $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$. By the assumptions, $\rho(T^{{\boldsymbol{s}}_{0}}x,T^{h({\boldsymbol{s}}_{0})}y)<\delta/3$, and by induction, $||h({\boldsymbol{s}}_{0})-{\boldsymbol{s}}_{0}||<2\delta/3$, so part (2) of Lemma 3.2 implies
@@
Since $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$, the argument above shows $(T^{{\boldsymbol{s}}_{0}}x)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]=(T^{{\boldsymbol{s}}_{0}+{\boldsymbol{r}}}y)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]$ for some $||{\boldsymbol{r}}||<\delta/3$. Equivalently, $x[\{{\boldsymbol{t}}_{0}\}]=(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]$. But also, the assumption, $\rho(T^{{\boldsymbol{t}}_{0}}x,T^{h({\boldsymbol{t}}_{0})}y)<\delta/3$ implies $x[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$ for some $||{\boldsymbol{r}}^{\prime}||<\delta/3$. So $(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$. As above, part (1) of Lemma 3.2 implies $h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}={\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}$, which yields the contradiction $32\delta/3=||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=||{\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}||<2\delta/$ 3.
@@
Assume $T$ is weakly expansive in the direction ${\mathcal{V}}$. Fix $\epsilon>0$ and find $\delta^{\prime}>0$ for weak expansiveness ((1) of Definition 2.10), satisfying $\delta^{\prime}<\Delta_{p}$. Let $x,y\in X$ and let $h=h_{x,y}$ be as in part (2) of Definition 2.10 satisfying: (i) $\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)<\delta^{\prime}/3$. Then by part (3) of Lemma 3.2 and Lemma 3.10 we have: (ii) $\rho(T^{\boldsymbol{t}}y,T^{h({\boldsymbol{t}})}y)=||h({\boldsymbol{t}})-{\boldsymbol{t}}||<2\delta^{\prime}/3$. Thus $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)\leq\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)+\rho(T^{h({\boldsymbol{t}})}y,T^{\boldsymbol{t}}y)<\delta^{\prime}/3+2\delta^{\prime}/3=\delta^{\prime}$, by (i) and (ii). It follows from weak expansiveness (part (1) of Definition 2.10) that $y=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{s}}_{0}||<\epsilon$. This implies $T$ is strongly expansive (for $\delta=\delta^{\prime}/3$) in the direction ${\mathcal{V}}$. ∎

### v2

@@ 2.6. Directional Expansiveness for $\mathbb{R}^{d}$
It turns out that the correct definition of direction ${\mathcal{V}}$ expansiveness for an $\mathbb{R}^{d}$ action $T$ is not simply the definition of expansiveness for $T$, restricted to ${\mathcal{V}}$, even though such a definition does work in some directional theories. The restriction of $T$ to ${\mathcal{V}}$, defined by $(T|_{\mathcal{V}})^{\boldsymbol{t}}:=T^{\boldsymbol{t}}$ for ${\boldsymbol{t}}\in{\mathcal{V}}$, is a ${\mathcal{V}}$ action, but can be made into an $\mathbb{R}^{e}$ action by a choice of a basis (see e.g., [3]). This does not affect questions of expansiveness.
@@ 3. Tiling dynamical systems
In this section we define finite local complexity (FLC) $\mathbb{R}^{2}$ tiling spaces and $\mathbb{R}^{2}$ tiling dynamical systems, and describe their basic properties (see e.g., [28] or [1]). Similar considerations hold for $\mathbb{R}^{d}$, $d>2$, but for simplicity we do not discuss them here.
@@ 3.1. Basic definitions
A tile is a subset $D\subseteq\mathbb{R}^{2}$ homeomorphic to a closed disk, with a Lebesgue measure zero boundary (e.g., a convex polygon). A tiling $x$ of $\mathbb{R}^{2}$ is a set of the tiles that cover $\mathbb{R}^{2}$, intersecting only on their boundaries. A tiling patch $x$ in $\mathbb{R}^{2}$ is a finite collection of the tiles whose union is a closed topological disc, and a tiling swatch $x$ is a collection of tiles whose union is closed and simply connected.
If $x$ is a tiling of $\mathbb{R}^{2}$, and $S\subseteq\mathbb{R}^{2}$, we let $x[S]$ be the smallest patch or swatch, the union of whose tiles contains $S$. Two tiles, patches, swatches or tilings are said to be equivalent if they are congruent via a translation. Sometimes geometrically equivalent tiles are “decorated” with colors or markings that make them inequivalent. We denote the translation of a tiling (or tile or patch or swatch) $x$ by ${{\boldsymbol{t}}}\in\mathbb{R}^{2}$ by $T^{{\boldsymbol{t}}}x:=\{D-{{\boldsymbol{t}}}:D\in x\}$. We call a set $p$ of inequivalent tiles in $\mathbb{R}^{2}$ a set of prototiles. A tiling (or path or swatch) $x$ by $p$ means $x$ each tile in $x$ is equivalent to a tile in $p$. We want to look at sets of all tilings by $p$, but without further restrictions, such a set is often too large.
@@
There is a well known compact metric topology on any FLC tiling spaces $X_{p}$ in which two tilings are close if, after a small translation, they agree on a large ball around the origin. Here is one version of such a metric $\rho$ (see [28], [34]).
Definition 3.2.
For each tile $D\in p$, let $\delta(D)=\inf\{||{\boldsymbol{r}}||:T^{\boldsymbol{r}}D\cap D=\emptyset\}$ and let $\Delta_{p}:=\min_{D\in p}\delta(D)$. For $x,y\in X_{p}$, $x\neq y$, let $R(x,y):=\sup\left\{r:(T^{{\boldsymbol{r}}_{x}}x)[B_{r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{r}}_{y}}y)[B_{r}({{\boldsymbol{0}}})]\text{ some }||{\boldsymbol{r}}_{x}||,||{\boldsymbol{r}}_{y}||<1/r\right\}$, Define the tiling metric $\rho(x,y):=\min(\Delta_{p},1/R(x,y))$.
A well known result of Rudolph [33] is that an FLC full tiling space $X_{p}$ is a compact metric space (see also [28], [34]). The following facts will be useful.
@@
    If $\rho(x,y)<\delta<\min(1/3,\Delta_{p})$ and $r=1/\delta$, then there exists a ${\boldsymbol{t}}\in\mathbb{R}^{d}$ with $||{\boldsymbol{t}}||<\delta$ so that $x[B_{(8/9)r}({{\boldsymbol{0}}})]=(T^{{\boldsymbol{t}}}y)[B_{(8/9)r}({{\boldsymbol{0}}})]$.
@@
For a full FLC tiling space $X_{p}$, we call any closed $T$-invariant subset $X\subseteq X_{p}$ an FLC tiling space (i.e., subspace) and call $(X,\mathbb{R}^{2},T)$ an FLC tiling dynamical system.
Let $X$ and $X^{\prime}$ be tiling spaces. For $r>0$, consider the sets of patches $X_{\{r\}}:=\{x[B_{r}({\boldsymbol{0}})]:x\in X\}$ and $X^{\prime}_{\{0\}}:=\{x^{\prime}[\{{\boldsymbol{0}}\}]:x^{\prime}\in X^{\prime}\}$.
Definition 3.5.
A map $C:X\to X^{\prime}$ is called local with local radius $r$ (see [28], [1]) if there is a map $c:X_{\{r\}}\to X^{\prime}_{\{0\}}$ so that $C(x)[\{{\boldsymbol{0}}\}]=c(x[B_{r}({\boldsymbol{0}})])$ for all $x\in X$.
Note that a local map is uniformly continuous, and satisfies $C(T^{\boldsymbol{t}}x)=T^{\boldsymbol{t}}C(x)$ for $x\in X$ and ${\boldsymbol{t}}\in\mathbb{R}^{2}$. A local map is a factor map if it is onto. If a local homeomorphism $C:X\to X^{\prime}$ has a local inverse, we say $X$ and $X^{\prime}$ are mutually locally derivable (MLD), and say $C$ is a MLD topological conjugacy. Local maps between FLC tiling spaces are analogous to sliding block codes between $\mathbb{Z}^{d}$ subshifts. In the subshift case, the Curtis/Lyndon/Hedlund Theorem says every topological conjugacy is a sliding block code. However, for FLC tiling dynamical systems there are topological conjugacies that are not MLD (see [26], [15], [1], and Theorem 6.11).
@@
For $j=0,1,\dots,4$, let ${{\boldsymbol{v}}}_{j}:=({\rm Re}(e^{2\pi ij/5}),{\rm Im}(e^{2\pi ij/5}))$, the $j$th $5$th root of unity, and let ${{\boldsymbol{v}}}^{\prime}_{j}:=(2/5){{\boldsymbol{v}}}_{j}$. We denote the prototile set by $p_{5}:=\{{{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:0\leq i<j\leq 4\}$, where ${{\boldsymbol{v}}}^{\prime}_{i}\wedge{{\boldsymbol{v}}}^{\prime}_{j}:=\{t_{0}{{\boldsymbol{v}}}^{\prime}_{i}+t_{1}{{\boldsymbol{v}}}^{\prime}_{j}:t_{0},t_{1}\in[0,1]\}$. Then $p_{5}$ consists of $10$ rhombic tiles with edge length $2/5$, five with acute angle $52\pi/$ 5 and five with acute angle $102\pi/$ 10. Figure 1 (a) shows ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{4}$ and ${{\boldsymbol{v}}}^{\prime}_{0}\wedge{{\boldsymbol{v}}}^{\prime}_{3}$. All other tiles in $p_{5}$ are rotations of one of these. We take local rule $p^{(2)}_{5}$ to be the requirement that all tiles meet edge to edge. The corresponding FLC full tiling space satisfies $X_{p_{5}}\neq\emptyset$ since it contains periodic tilings (see Figure 1, (b)).
@@ 3.2. Expansiveness and directional expansiveness for tiling dynamical systems
Given $\epsilon>0$, let $\delta<\min\left(\epsilon,1/3,\Delta_{p}/2\right)$, $r=1/\delta$. Let $B_{0}=B_{8r/9}({{\boldsymbol{0}}})$. For a sequence ${\boldsymbol{t}}_{n}\in\mathbb{R}^{2}$, $n=1,2,\dots$, let $B_{n}:=B_{8r/9}(-{\boldsymbol{t}}_{n})$. Choose the ${\boldsymbol{t}}_{n}$ so that $\cup_{n\geq 1}B_{n}=\mathbb{R}^{2}$, and so that $I_{n}:=B_{n+1}\cap B_{n}\neq\emptyset$. Now suppose $x,y\in X_{p}$ satisfy $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)<\delta$ for all ${\boldsymbol{t}}\in\mathbb{R}^{2}$. Then for each $n=1,2,\dots$, $\rho(T^{{\boldsymbol{t}}_{n}}x,T^{{\boldsymbol{t}}_{n}}y)<\delta$, so that by (2) of Lemma 3.3, there exists $||{\boldsymbol{r}}_{n}||<\delta$ so that $(T^{{\boldsymbol{t}}_{n}}x)[B_{0}]=(T^{{\boldsymbol{t}}_{n}+{\boldsymbol{r}}_{n}}y)[B_{0}]$. Equivalently, $x[B_{n}]=(T^{{\boldsymbol{r}}_{n}}y)[B_{n}]$. We claim ${\boldsymbol{r}}_{n}={\boldsymbol{r}}_{1}$ for all $n\in\mathbb{N}$. Thus $x[B_{n}]=(T^{{\boldsymbol{r}}_{1}}y)[B_{n}]$ for all $n$, and since the $\mathbb{R}^{2}=\cup_{n\geq 1}B_{n}$, $y=T^{{\boldsymbol{r}}_{1}}x$. By (3) of Lemma 3.3, $||{\boldsymbol{r}}_{1}||=\rho(x,y)<\delta<\epsilon$. To prove the claim, we prove ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$ for all $n$. Let $C_{n}:=B_{n}\cap B_{n+1}$. Then $(T^{{\boldsymbol{r}}_{n}}y)[C_{n}]=x[C_{n}]=(T^{{\boldsymbol{r}}_{n+1}}y)[C_{n}]$. Since $||{\boldsymbol{r}}_{n+1}-{\boldsymbol{r}}_{n}||<2\delta<\Delta_{p}$, (1) of Lemma 3.3 implies ${\boldsymbol{r}}_{n+1}={\boldsymbol{r}}_{n}$. ∎
@@
For the induction step, we assume that, for some $k\geq 1$, no ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<rk$ has $||h({\boldsymbol{t}})-{\boldsymbol{t}}||\geq 2\delta/3$, and show the same for ${\boldsymbol{t}}\in{\mathcal{V}}$ with $||{\boldsymbol{t}}||<(k+1)r$. Suppose, to the contrary, that some such ${\boldsymbol{t}}={\boldsymbol{t}}_{0}$ satisfies $||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=2\delta/3$. Let ${\boldsymbol{s}}_{0}=kt/(k+1)$ so that ${\boldsymbol{s}}_{0}\in{\mathcal{V}}$, $||{\boldsymbol{s}}_{0}||<kr$ and $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$. By the assumptions, $\rho(T^{{\boldsymbol{s}}_{0}}x,T^{h({\boldsymbol{s}}_{0})}y)<\delta/3$, and by induction, $||h({\boldsymbol{s}}_{0})-{\boldsymbol{s}}_{0}||<2\delta/3$, so part (2) of Lemma 3.3 implies
@@
Since $||{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}||<r$, the argument above shows $(T^{{\boldsymbol{s}}_{0}}x)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]=(T^{{\boldsymbol{s}}_{0}+{\boldsymbol{r}}}y)[{\boldsymbol{t}}_{0}-{\boldsymbol{s}}_{0}]$ for some $||{\boldsymbol{r}}||<\delta/3$. Equivalently, $x[\{{\boldsymbol{t}}_{0}\}]=(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]$. But also, the assumption, $\rho(T^{{\boldsymbol{t}}_{0}}x,T^{h({\boldsymbol{t}}_{0})}y)<\delta/3$ implies $x[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$ for some $||{\boldsymbol{r}}^{\prime}||<\delta/3$. So $(T^{\boldsymbol{r}}y)[\{{\boldsymbol{t}}_{0}\}]=(T^{h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}+{\boldsymbol{r}}^{\prime}}y)[\{{\boldsymbol{t}}_{0}\}]$. As above, part (1) of Lemma 3.3 implies $h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}={\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}$, which yields the contradiction $32\delta/3=||h({\boldsymbol{t}}_{0})-{\boldsymbol{t}}_{0}||=||{\boldsymbol{r}}^{\prime}-{\boldsymbol{r}}||<2\delta/$ 3. ∎
@@
Assume that $T$ is weakly expansive in the direction ${\mathcal{V}}$. Fix $\epsilon>0$ and find $\delta^{\prime}>0$ for weak expansiveness ((1) of Definition 2.9), satisfying $\delta^{\prime}<\Delta_{p}$. Let $x,y\in X$ and let $h=h_{x,y}$ be as in part (2) of Definition 2.9 satisfying: (i) $\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)<\delta^{\prime}/3$. Then by part (3) of Lemma 3.3 and Lemma 3.11 we have: (ii) $\rho(T^{\boldsymbol{t}}y,T^{h({\boldsymbol{t}})}y)=||h({\boldsymbol{t}})-{\boldsymbol{t}}||<2\delta^{\prime}/3$. Thus $\rho(T^{\boldsymbol{t}}x,T^{\boldsymbol{t}}y)\leq\rho(T^{\boldsymbol{t}}x,T^{h({\boldsymbol{t}})}y)+\rho(T^{h({\boldsymbol{t}})}y,T^{\boldsymbol{t}}y)<\delta^{\prime}/3+2\delta^{\prime}/3=\delta^{\prime}$, by (i) and (ii). It follows from weak expansiveness (part (1) of Definition 2.9) that $y=T^{{\boldsymbol{s}}_{0}}x$ for some ${\boldsymbol{s}}_{0}\in\mathbb{R}^{2}$ with $||{\boldsymbol{s}}_{0}||<\epsilon$. This implies $T$ is strongly expansive (for $\delta=\delta^{\prime}/3$) in the direction ${\mathcal{V}}$. ∎